# Continuous Approximations of Projected Dynamical Systems via Control Barrier Functions

### v1

@@ Abstract
Projected Dynamical Systems (PDSs) form a class of discontinuous constrained dynamical systems, and have been used widely to solve optimization problems and variational inequalities. Recently, they have also gained significant attention for control purposes, such as high-performance integrators, saturated control and feedback optimization. In this work, we establish that locally Lipschitz continuous dynamics, involving Control Barrier Functions (CBFs), namely CBF-based dynamics, approximate PDSs. Specifically, we prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDSs, as a CBF-parameter is taken to infinity. Towards this, we also prove that CBF-based dynamics are perturbations of PDSs, with quantitative bounds on the perturbation. Our results pave the way to implement discontinuous PDS-based controllers in a continuous fashion, employing CBFs. Moreover, they can be employed to numerically simulate PDSs, overcoming disadvantages of existing discretization schemes, such as computing projections to possibly non-convex sets. Finally, this bridge between CBFs and PDSs may yield other potential benefits, including novel insights on stability.
@@ I Introduction
In this work, we reveal an intimate relationship between PDSs and CBFs: we show that PDSs can be approximated by (locally Lipschitz) continuous dynamics, employing CBFs. Specifically, we present results formally establishing the fact that system interconnections containing CBF-based continuous dynamic controllers (hereby termed CBF-based dynamics) approximate interconnections with discontinuous PDSs (hereby termed PDS-based dynamics). Our results pave the way for implementing the PDS-based feedback controllers mentioned above in a continuous fashion, employing CBFs, with all the advantages that may come with continuity, such as increased robustness and easier sampled-data implementations.
Moreover, our work has implications for purposes other than PDS-based control as well. First, one may employ locally Lipschitz CBF-based dynamics to numerically simulate discontinuous PDSs. While simulating Lipschitz CBF-based dynamics is computationally easy, standard PDS discretization schemes (see e.g. [1, 18]) involve computing a projection to $\mathcal{S}$, which is computationally expensive, especially when $\mathcal{S}$ is non-convex. This may be particularly useful in online scenarios, such as model predictive control and online optimization. Further, our results build bridges between PDS theory and CBF theory. E.g., in our preliminary work [19], we demonstrated how, in simple scenarios, one may infer stability of CBF-based control from the associated PDS, addressing the notorious undesired equilibrium problem of CBFs (cf. [16]) [^2] Note, however, that, in our scenario, there is no restriction on modifying the dynamics, as it is the controller’s dynamics that is modified, whereas in the standard CBF-literature there are restrictions on modifying the dynamics. In standard CBF literature, the goal is designing a controller $u(x)$, such that trajectories of a system $\dot{x}=f(x)+g(x)u$ remain in a set $\mathcal{S}$. In contrast, we focus on freely shaping the dynamics of a dynamic controller; this reduces to the above scenario, if the dynamics is simplified to $\dot{x}=f(x)+u$..
@@
We prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDS-based dynamics, as the tunable parameter $\alpha$ in the CBF constraint is taken to $\infty$. Thus, as $\alpha$ becomes larger, CBF-based dynamics approximate more accurately PDS-based dynamics. This supports the use of CBF-based dynamics to implement PDS-based dynamics, as, for large $\alpha$, their trajectories are arbitrarily close. Towards proving convergence of trajectories, we also show that CBF-based dynamics are perturbations of PDSs, in the sense that the CBF-based vector field belongs in a set of perturbations of the set-valued map of the differential inclusion that describes the PDS. We obtain quantitative bounds on the perturbation, that depend on and vanish with $\alpha$. Finally, we demonstrate our results on a feedback-optimization scenario, showing how PDS-based controllers can be implemented in a continuous fashion, employing CBFs.
@@
Our preliminary work [19] explored the relationship between PDSs and CBFs. Focusing on autonomous dynamics, it proved that CBF-based dynamics are perturbed versions of PDSs [19, Thm. III.1]. Here, we extend [19, Thm. III.1] in two ways: a) we consider the more general non-autonomous case of interconnections of systems with dynamics that are either PDS-like or CBF-based, and b) we provide bounds on the perturbed set-valued map, establishing local boundedness thereof, which is necessary for proving convergence of trajectories. What is more, convergence of trajectories - the present article’s main theme - is not studied in [19].
The works [20, 21, 22] also employ CBF-based dynamics instead of PDS-based ones for constrained optimization problems, variational inequalities and feedback optimization. Mainly, they focus on the relationship between the equilibria of CBF-based and PDS-based dynamics, and provide asymptotic stability and contractivity results on CBF-based dynamics. Further, in [20], it is proven that the vector-fields of PDSs and CBF-based dynamics coincide at $\alpha=\infty$. However, neither convergence of trajectories nor the fact that CBF-based dynamics are perturbations of PDS-based ones, with quantitative bounds on the perturbation, depending on $\alpha$, is established. Overall, our work can be viewed as complementary to [20, 21, 22].
Finally, [10] proposed a type of continuous approximations of PDSs, based on antiwindup control, called Anti-Windup Approximations (AWA). Some steps on establishing convergence of trajectories in our work are inspired by [10] (e.g. proving that CBF-based dynamics are perturbations of PDS-based ones). Nonetheless, as CBF-based dynamics is different than AWA, proving the corresponding results is also significantly different. Further, trajectories of AWA, contrary to CBF-based dynamics, stay in an inflation of $\mathcal{S}$. If one needs to force AWA to stay in $\mathcal{S}$, modifications have to be made, which may result in undesired behaviors, such as discarding useful equilibria (e.g. extrema of feedback optimization) or shrinking regions of attraction. In addition, AWA involve calculating projections to $\mathcal{S}$, which can be computationally expensive. On the other hand, when $\mathcal{S}$ represents limitations of the physical system (e.g., input saturation), and said projections are physically enforced by the system, implementation of AWA does not even require knowledge of $\mathcal{S}$, in contrast to CBF-based dynamics. Apart from the above, both our work and [10] establish interesting, previously unknown, connections between PDSs and other popular classes of dynamics, and might even suggest connections between antiwindup control and CBFs.
@@ II-A Notation
Denote the set of positive-definite symmetric matrices in ${}^{n\times n}$ by $\mathbb{S}_{+}^{n}$. Given $P\in\mathbb{S}_{+}^{n}$, denote its minimum and maximum eigenvalues by $\underline{\lambda}(P)$ and $\overline{\lambda}(P)$, respectively. Given $P\in\mathbb{S}_{+}^{n}$ and $x\in{}^{n}$, denote $\|x\|_{P}=\sqrt{x^{\top}Px}$. Given a function $f:{}^{m}\times{}^{n}\to{}^{n}$ and a differentiable function $h:{}^{n}\to\real$, denote $\mathcal{L}_{f}h(z,x):=\nabla h^{\top}(x)\cdot f(z,x)$, where $(z,x)\in{}^{m}\times{}^{n}$. A continuous function $\gamma:[0,\infty)\to\real$ is said to belong to $\mathcal{K}_{\infty}$, if $\gamma(0)=0$, $\gamma$ is strictly increasing, and $\lim_{a\to\infty}\gamma(a)=+\infty$. Finally, denote the closed unit ball of appropriate dimensions by $\mathbb{B}$.
@@ II-B Variational Analysis
where $h:{}^{n}\to\real$ is continuously differentiable, $x\mapsto\nabla h(x)$ is locally Lipschitz, and $\nabla h(x)\neq 0$ for all $x\in\partial\mathcal{S}$, are prox-regular and the expressions for their tangent and normal cones are obtained as follows (see [4, Example 6.8]), for any $x\in\mathcal{S}$:
@@ II-C Differential Inclusions
where $F:{}^{n}\rightrightarrows{}^{n}$ and we have omitted the dependence of $\xi$ on $t$, for conciseness. When it is clear from the context that the DI’s constraint set is ${}^{n}$, we might omit it. Given an initial condition $x_{0}\in\mathcal{S}$ and a $T\in[0,\infty)$, a function $\phi:[0,T]:\to{}^{n}$ is a (Carathéodory) solution to DI (3), if it is absolutely continuous, $\phi(0)=x_{0}$, $\phi(t)\in\mathcal{S}$ and $\dot{\phi}(t)\in F(\phi(t))$ for almost all $t\in[0,T]$. We adopt the same definition for solutions of differential equations, as well.
@@ III Background on Projected Dynamical Systems and Control Barrier Functions
$\mathcal{S}$ and $h:{}^{n}\to\real$ satisfy the following:
@@
Items 2 and 3, as well as item 1 without the compactness assumption, are standard in the literature of CBFs. Compactness of $\mathcal{S}$ is needed for several bounds in the proof of Prop. IV.1. Nonetheless, in our context, it is not a restrictive assumption, since, later on, $\mathcal{S}$ is the set where the controller’s state is confined, which is not expected to be unbounded in practice. In fact, as $\mathcal{S}$ can be taken arbitrarily large, our results are semiglobal. Further, item 4 holds a-priori if $h$ is a real-analytic function, by the Łojasiewicz inequality (see e.g. [26]). Also, as aforementioned, items 2 and 3 imply that $\mathcal{S}$ is prox-regular and its tangent and normal cones are expressed as in (2). Finally, the following standard assumption on $f$ is considered:
@@ III-A Projected Dynamical Systems
Systems of the form (5) are called Projected Dynamical Systems (PDSs) [1]. When $\xi\in\mathrm{Int}(\mathcal{S})$, we have $T_{\mathcal{S}}(\xi)={}^{n}$, and thus $\Pi^{P}_{\mathcal{S}}\Big{(}\xi,f(\xi)\Big{)}=f(\xi)$, i.e. the PDS (5) evolves according to the unconstrained dynamics (4). However, when $\xi\in\partial\mathcal{S}$, a vector $\mu\in T_{\mathcal{S}}(\xi)$ is chosen (minimizing $\|\mu-f(\xi)\|_{P}^{2}$), so that the trajectory remains in $\mathcal{S}$. Finally, when $\mathcal{S}$ is prox-regular, the right-hand side of (5) is a singleton and (5) becomes a discontinuous ODE.
@@
The interpretation of (6) is: when $\xi(t)$ is about to leave $\mathcal{S}$ (i.e. $\xi(t)\in\partial\mathcal{S}$), an element $\eta$ from the truncated normal cone $P^{-1}N_{\mathcal{S}}(\xi)\cap d(\xi)\mathbb{B}$ is chosen, such that $f(\xi)-\eta$ points inwards $\mathcal{S}$, thereby keeping $\xi(t)\in\mathcal{S}$. The following is a collection of results on solutions of PDSs from [4, 10]:
@@
    Let $\mathcal{S}$ be Clarke regular and $f$ be continuous. Consider $d:{}^{n}\to[0,\infty]$, such that $d(x)\geq\sqrt{\tfrac{\overline{\lambda}(P)}{\underline{\lambda}(P)}}\|f(x)\|$ for all $x\in\mathcal{S}$. Then, for every initial condition in $\mathcal{S}$, (5) and (6) admit the same solutions.
@@
where $f:{}^{m}\times{}^{n}\to{}^{n}$ $g:{}^{m}\times{}^{n}\to{}^{m}$. We interpret the above setup as: $\zeta(t)$ is the state of some system controlled by a dynamic controller $\xi(t)$, with PDS dynamics. Such interconnections have been proposed e.g. for saturated control [8, 9] and feedback optimization [4] [^4] A generalization of (7), termed extended PDSs, has also been employed in [5, 6] for high-performance integral control and [7] for passivity-based control. Extension of our results to ePDSs is considered future work.. The above interconnection can be written as:
@@ III-B Control-Barrier-Function-based dynamics
Given $f:{}^{n}\to{}^{n}$, a tunable parameter $\alpha>0$ and $P\in\mathbb{S}_{+}^{n}$, consider
@@
We refer to (9) as the CBF-based dynamics. The term comes from the CBF literature [13], which refers to $h$ as the control barrier function (CBF). Under Assumption 1, the Quadratic Program (QP) above has a unique solution, for any $\xi$, which can be written in closed form as follows (see [17, Thm. 1]):
@@
We call $f^{P}_{\mathrm{cbf},\alpha}$ the CBF-based vector field. In contrast to PDSs (5), here the nominal dynamics is modified even when $\xi\in\mathrm{Int}\mathcal{S}$. The nominal vector field $f$ is modified whenever $\mathcal{L}_{f}h(\xi)+\alpha h(\xi)\leq 0$, i.e. whenever the value of $h$ decreases along the trajectories of the nominal system (4) faster than a state-dependent threshold $-\alpha h(\xi)$. The larger $\alpha$ is, the less invasive the modification becomes, as it allows for even more negative decreases $\mathcal{L}_{f}h(\xi)$. Typically, this allows trajectories to approach closer to the boundary, before the nominal vector-field is modified.
Vector fields like (9) arise in control systems of the form $\dot{x}=f(x)+u(x)$, where the controller $u(x)$ is designed via CBF-based methods. From standard CBF theory [13], it follows that trajectories of (9) starting in $\mathcal{S}$ stay in $\mathcal{S}$. Under Assumptions 1 and 2, $f^{P}_{\mathrm{cbf},\alpha}$ is locally Lipschitz, implying existence and uniqueness of solutions of (9).
@@
where $f:{}^{m}\times{}^{n}\to{}^{n}$ and $g:{}^{m}\times{}^{n}\to{}^{m}$. The article’s main objective is to show that the feedback interconnection with a CBF-based controller (10) is a legitimate continuous approximation of the feedback interconnection with a PDS-based controller (7), in the sense of uniform convergence of trajectories. Briefly, as $\alpha\to\infty$, trajectories of (10) become identical to trajectories of (7). Towards this, we also derive another highly relevant result: that (10) is a perturbation of (7), with quantitative bounds on the perturbation.
@@
Note that all our results relating (trajectories of) (10) with (trajectories of) (7), also hold for (9) and (5), as (9) and (5) are special cases of (10) and (7), respectively. normal-$\blacksquare$

### v3

@@ Abstract
Projected Dynamical Systems (PDSs) form a class of discontinuous constrained dynamical systems, and have been used widely to solve optimization problems and variational inequalities. Recently, they have also gained significant attention for control purposes, such as high-performance integrators, saturated control and feedback optimization. In this work, we establish that locally Lipschitz continuous dynamics, involving Control Barrier Functions (CBFs), namely CBF-based dynamics, approximate PDSs. Specifically, we prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDSs, as a CBF-parameter approaches infinity. Towards this, we also prove that CBF-based dynamics are perturbations of PDSs, with quantitative bounds on the perturbation. Our results pave the way to implement discontinuous PDS-based controllers in a continuous fashion, employing CBFs. We demonstrate this on numerical examples on feedback optimization and synchronverter control. Moreover, our results can be employed to numerically simulate PDSs, overcoming disadvantages of existing discretization schemes, such as computing projections to possibly non-convex sets. Finally, this bridge between CBFs and PDSs may yield other potential benefits, including novel insights on stability.
@@ I Introduction
In this work, we reveal an intimate relationship between PDSs and CBFs: we show that PDSs can be approximated by (locally Lipschitz) continuous dynamics, employing CBFs. Specifically, we present results formally establishing the fact that system interconnections containing CBF-based continuous dynamics (hereby termed CBF-based dynamics) approximate interconnections with discontinuous PDSs (hereby termed PDS-based dynamics). Our results pave the way for implementing the PDS-based feedback controllers mentioned above in a continuous fashion, employing CBFs, with all the advantages that may come with continuity, such as increased robustness and easier sampled-data implementations.
Moreover, our work has implications for purposes other than PDS-based control as well. First, one may employ locally Lipschitz CBF-based dynamics to numerically simulate discontinuous PDSs, e.g. for constrained optimization. While simulating Lipschitz CBF-based dynamics is computationally easy, standard PDS discretization schemes (see e.g., [1, 18]) involve computing a projection to $\mathcal{S}$, which is computationally expensive, especially when $\mathcal{S}$ is non-convex. This may be particularly useful for online optimization (e.g., model predictive control). Further, our results build bridges between PDS theory and CBF theory. E.g., in our preliminary work [19], we showed how, in simple scenarios, one may infer stability of safety-critical CBF-based control from the associated PDS, addressing the notorious undesired equilibrium problem of CBFs (cf. [16]).
@@
We prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDS-based dynamics, as the tunable parameter $\alpha$ in the CBF constraint approaches $\infty$. Thus, as $\alpha$ becomes larger, CBF-based dynamics approximate more accurately PDS-based dynamics. This supports the use of CBF-based dynamics to implement PDS-based dynamics, as, for large $\alpha$, their trajectories are arbitrarily close. Towards proving convergence of trajectories, we also show that CBF-based dynamics are perturbations of PDSs, in the sense that the CBF-based vector field belongs in a set of perturbations of the set-valued map of the differential inclusion that describes the PDS. We obtain quantitative bounds on the perturbation, that depend on and vanish with $\alpha$. Finally, we demonstrate our results on examples on feedback optimization and synchronverter control, implementing PDS-based controllers from the literature in a continuous fashion, via CBFs.
@@
Our preliminary work [19] explored the relationship between PDSs and CBFs. However, [19] did not study convergence of trajectories, which is the present article’s main theme. [^2] As such, Cors. IV.1, IV.2 and IV.3, Thms. IV.1 and IV.2, and Prop. VIII.1 appear here for the first time. Further, [19], focusing on autonomous dynamics, proved that CBF-based dynamics are perturbed versions of PDSs [19, Thm. III.1]. Here, we extend [19, Thm. III.1] in two ways: a) we consider the more general non-autonomous case of interconnections of systems with dynamics that are either PDS-like or CBF-based, and b) we provide bounds on the perturbed set-valued map, establishing local boundedness thereof, which is necessary for proving convergence of trajectories.
The works [20, 21, 22] also employ CBF-based dynamics instead of PDS-based ones for constrained optimization problems, variational inequalities and feedback optimization. Mainly, they focus on the relationship between the equilibria of CBF-based and PDS-based dynamics, and provide asymptotic stability and contractivity results on CBF-based dynamics. Further, in [20], it is proven that the right-hand sides of the differential equations of PDSs and CBF-based dynamics coincide at $\alpha=\infty$. However, neither convergence of trajectories nor the fact that CBF-based dynamics are perturbations of PDS-based ones, with quantitative bounds on the perturbation, depending on $\alpha$, is established. Overall, our work can be viewed as complementary to [20, 21, 22].
Finally, [10] proposed a type of continuous approximations of PDSs, based on antiwindup control, called Anti-Windup Approximations (AWA). Some steps on establishing convergence of trajectories in our work are inspired by [10] (e.g., proving that CBF-based dynamics are perturbations of PDS-based ones). Nonetheless, as CBF-based dynamics are different than AWA, proving the corresponding results is also significantly different. Further, trajectories of AWA, contrary to CBF-based dynamics, stay in an inflation of $\mathcal{S}$. If one needs to force AWA to stay in $\mathcal{S}$, modifications have to be made, which may result in undesired behaviors, such as discarding useful equilibria (e.g., extrema of feedback optimization) or shrinking regions of attraction. In addition, AWA involve calculating projections to $\mathcal{S}$, which can be computationally expensive. On the other hand, when $\mathcal{S}$ represents limitations of the physical system (e.g., input saturation), and said projections are physically enforced by the system, implementation of AWA does not even require knowledge of $\mathcal{S}$, in contrast to CBF-based dynamics. Apart from the above, both our work and [10] establish interesting, previously unknown, connections between PDSs and other popular classes of dynamics, and might even suggest connections between antiwindup control and CBFs.
@@ II-A Notation
Denote the set of positive-definite symmetric matrices in n×n by $\mathbb{S}_{+}^{n}$. Given $P\in\mathbb{S}_{+}^{n}$, denote its minimum and maximum eigenvalues by $\underline{\lambda}(P)$ and $\overline{\lambda}(P)$, respectively. Given $P\in\mathbb{S}_{+}^{n}$ and $x\in^{n}$, denote $\|x\|_{P}=\sqrt{x^{\top}Px}$. Given a function $f:^{m}\times^{n}\to^{n}$ and a differentiable function $h:^{n}\to$, denote $\mathcal{L}_{f}h(z,x):=\nabla h^{\top}(x)\cdot f(z,x)$, where $(z,x)\in^{m}\times^{n}$. A continuous function $\gamma:[0,\infty)\to$ is said to belong to $\mathcal{K}_{\infty}$, if $\gamma(0)=0$, $\gamma$ is strictly increasing, and $\lim_{a\to\infty}\gamma(a)=+\infty$. Finally, denote the closed unit ball of appropriate dimensions by $\mathbb{B}$.
@@ II-B Variational Analysis
where $h:^{n}\to$ is continuously differentiable, $x\mapsto\nabla h(x)$ is locally Lipschitz, and $\nabla h(x)\neq 0$ for all $x\in\partial\mathcal{S}$, are prox-regular and the expressions for their tangent and normal cones are obtained as follows (see [4, Example 6.8]), for any $x\in\mathcal{S}$:
@@ II-C Differential Inclusions
where $F:^{n}\rightrightarrows^{n}$ and we have omitted the dependence of $\xi$ on $t$, for conciseness. When it is clear from the context that the DI’s constraint set is n, we might omit it. Given an initial condition $x_{0}\in\mathcal{S}$ and a $T\in[0,\infty)$, a function $\phi:[0,T]:\to^{n}$ is a (Carathéodory) solution to DI (3), if it is absolutely continuous, $\phi(0)=x_{0}$, $\phi(t)\in\mathcal{S}$ and $\dot{\phi}(t)\in F(\phi(t))$ for almost all $t\in[0,T]$. We adopt the same definition for solutions of differential equations, as well.
@@ III Background on Projected Dynamical Systems and Control Barrier Functions
$\mathcal{S}$ and $h:^{n}\to$ satisfy the following:
@@
Items 2, 3 and item 1 without the compactness assumption, are standard in the literature of CBFs (see e.g., [13] or [17]). Compactness of $\mathcal{S}$ is needed for several bounds in the proof of Prop. IV.1. Nonetheless, in many related applications, such as PDS-based control, this is not a restrictive assumption. In fact, as $\mathcal{S}$ can be taken arbitrarily large, our results are semiglobal. Further, item 4 holds a-priori if $h$ is a real-analytic function, by the Łojasiewicz inequality (see, e.g.,[26]). Also, as aforementioned, items 2 and 3 imply that $\mathcal{S}$ is prox-regular and its tangent and normal cones are expressed as in (2). Finally, the following standard assumption on $f$ is considered:
@@ III-A Projected Dynamical Systems
Systems of the form (5) are called Projected Dynamical Systems (PDSs) [1]. When $\xi\in\mathrm{Int}(\mathcal{S})$, we have $T_{\mathcal{S}}(\xi)=^{n}$, and thus $\Pi^{P}_{\mathcal{S}}\Big{(}\xi,f(\xi)\Big{)}=f(\xi)$, i.e., the PDS (5) evolves according to the unconstrained dynamics (4). However, when $\xi\in\partial\mathcal{S}$, a vector $\mu\in T_{\mathcal{S}}(\xi)$ is chosen (minimizing $\|\mu-f(\xi)\|_{P}^{2}$), so that the trajectory remains in $\mathcal{S}$. Finally, when $\mathcal{S}$ is prox-regular, the right-hand side of (5) is a singleton and (5) becomes a discontinuous ODE.
@@
The interpretation of (6) is: when $\xi(t)$ is about to leave $\mathcal{S}$ (i.e., $\xi(t)\in\partial\mathcal{S}$), an element $\eta$ from the truncated normal cone $P^{-1}N_{\mathcal{S}}(\xi)\cap d(\xi)\mathbb{B}$ is chosen, such that $f(\xi)-\eta$ points inwards $\mathcal{S}$, thereby keeping $\xi(t)\in\mathcal{S}$. The following is a collection of results on solutions of PDSs from [4, 10]:
@@
    Let $\mathcal{S}$ be Clarke regular and $f$ be continuous. Consider $d:^{n}\to[0,\infty]$, such that $d(x)\geq\boldsymbol{s}\sqrt{\tfrac{\overline{\lambda}(P)}{\underline{\lambda}(P)}}\|f(x)\|$ for all $x\in\mathcal{S}$. Then, for every initial condition in $\mathcal{S}$, (5) and (6) admit the same solutions.
@@
where $f:^{m}\times^{n}\to^{n}$ $g:^{m}\times^{n}\to^{m}$. The above setup can have different interpretations. For example, $\zeta(t)$ may be the state of some system controlled by a dynamic controller $\xi(t)$, with PDS dynamics. Such interconnections have been proposed e.g., for saturated control [8, 9] and feedback optimization [4]. [^4] A generalization of (7), termed extended PDSs, has also been employed in [5, 6] for high-performance integral control and [7] for passivity-based control. Extension of our results to ePDSs is considered future work. Moreover, one can disregard the $\zeta$-part (i.e., $g(\zeta,\xi)=0$) in (7), and obtain a standard PDS. Thus, our work also applies to cases such as constrained optimization [1, 2, 3, 4]. The above interconnection can be written as:
@@ III-B Control-Barrier-Function-based dynamics
Given $f:^{n}\to^{n}$, a parameter $\alpha>0$ and $P\in\mathbb{S}_{+}^{n}$, consider
@@
We refer to (9) as the CBF-based dynamics as $h$ is referred to as the control barrier function (CBF) in the CBF literature [13]. Under Assumption 1, the QP above has a unique solution, for any $\xi$, which can be written in closed form as follows (see [17, Thm. 1]):
@@
We call $f^{P}_{\mathrm{cbf},\alpha}$ the CBF-based vector field. In contrast to PDSs (5), the nominal dynamics is modified even when $\xi\in\mathrm{Int}\mathcal{S}$; $f$ is modified whenever $\mathcal{L}_{f}h(\xi)+\alpha h(\xi)\leq 0$, i.e., whenever $h$ decreases along the nominal system’s trajectories (4) faster than a state-dependent threshold $-\alpha h(\xi)$. The larger $\alpha$ is, the less invasive the modification is, as it allows for even more negative decreases $\mathcal{L}_{f}h(\xi)$. Typically, this allows trajectories to approach closer to the boundary, before the nominal vector-field is modified.
Vector fields like (9) arise in control systems of the form $\dot{x}=f(x)+u(x)$, where the controller $u(x)$ is designed via CBF-based methods. From standard CBF theory [13], trajectories of (9) starting in $\mathcal{S}$ stay in $\mathcal{S}$. Under Assumptions 1 and 2, $f^{P}_{\mathrm{cbf},\alpha}$ is locally Lipschitz, implying existence and uniqueness of solutions of (9).
@@
where $f:^{m}\times^{n}\to^{n}$ and $g:^{m}\times^{n}\to^{m}$. The article’s main objective is to show that the feedback interconnection with a CBF-based controller (10) is a continuous approximation of the feedback interconnection with a PDS-based controller (7), in the sense of uniform convergence of trajectories. Briefly, as $\alpha\to\infty$, trajectories of (10) become identical to trajectories of (7). Towards this, we also derive another highly relevant result: that (10) is a perturbation of (7), with quantitative bounds on the perturbation.
@@
Note that all our results relating (trajectories of) (10) with (trajectories of) (7), also hold for (9) and (5), as (9) and (5) are special cases of (10) and (7), respectively. $\blacksquare$